# Imaging the positron annihilation line with 20-year INTEGRAL/SPI observations

### v2

@@ 4.3.2 Other regions
The reconstructed image shows a blob-like structure at high Galactic latitude around ($l$, $b$) = (70, 40). When defining the region as $l=60^{\circ}$ to $90^{\circ}$, $b=30^{\circ}$ to $60^{\circ}$, we measure the flux of this region to be $(7.5\pm_{\mathrm{stat}}2.4\pm_{\mathrm{sys}}3.4)\times 10^{-5}$ ph cm-2 s-1. None of the background-only bootstrap samples reproduce such a high flux in this region. Since no known bright high-energy source has been identified in this region, it could be an interesting source emitting 511 keV line predominantly, e.g., a locally dense region aligned with stellar streams, Galactic winds, or transient events such as binary neutron star mergers (Fuller et al. 2019) and calcium-rich supernovae (Perets 2014), or nearby bubbles (Siegert et al. 2024).
@@
Although it is statistically significant, we need to be careful about artifacts of the image deconvolution. As seen at the bottom of Figure 8, the exposure is not uniform in high-latitude regions. Furthermore, when splitting the full dataset into four independent periods, we find that this structure appears obviously in the image of the last period (years 2017–2023, see Appendix F and Figure 18). From January 2022 to October 2022, SPI monitored the blazar Mkn 501, located at ($l$, $b$) = (63.6, 38.9) and apart by $\sim 6^{\circ}$ from the center of the blob-like structure. This feature may arise from systematic uncertainties in the SPI response function at the boundaries of its field of view rather than being a real astronomical source. Additionally, we observe a similar structure when performing image reconstruction in an adjacent energy band (526–532 keV) with a flux of $\sim 4\times 10^{-5}$ ph cm-2 s-1.
@@
At ($l$, $b$) = ($260$, 60-60) degrees, another structure can be seen, annotated as “South Blob” in Figure 8. However, the non-uniform exposure could also affect this structure. To conclude the existence of these structures, uniform exposure is crucial to minimize such systematic uncertainties in image reconstruction for future 511 keV observations. In any case, this southern emission feature is too far away from the Magellanic Clouds. Interestingly, Siegert et al. (2020) saw a similar emission feature at high negative latitudes with the COSI balloon, which they also attributed to image deconvolution artifacts.
@@
Other emission features in Figure 8, such as around the Cepheus region ($l\sim 120^{\circ}$) or at the position of M82 (in particular deep exposure of the type Ia supernova SN2014J, e.g, Diehl et al. (2014); $(l,b)\sim(140^{\circ},40^{\circ})$) appear as reasonable positron annihilation regions, but are not significant. Local features, such as Loop I or the North Polar Spur, are also not aligned with any of these serendipitous emission features. Strong individual hotspots from the filamentary structure of the Local Bubble (Siegert et al. 2024) may explain some residual emission, however not at the flux levels quoted here.
@@
Table 4: Measured fluxes and significance of the 511 keV line emission from individual regions.
@@
 [^4] The chance probabilities are derived from background-only bootstrap samples.
@@ 5.1 Comparison with previous studies
For example, Weidenspointner et al. (2008) suggested that low-mass X-ray binaries with hard X-ray emissions (hard LMXBs) explain the 511 keV emission in the disk region. They reported a flux ratio of $\sim 2$ between negative and positive longitude regions, and hard LMXBs show a similar asymmetric distribution along the Galactic plane. To revisit this hypothesis using our 20-year dataset, we examined the spatial distribution of LMXBs using the Swift-BAT 157-month catalog (Lien et al. 2025). The catalog shows 71 LMXBs in the eastern hemisphere ($l<0^{\circ}$) versus 46 in the west ($l>0^{\circ}$), confirming an asymmetric distribution similar to previous findings (see Figure 10). However, the interpretation depends critically on the assumed positron production efficiency. When considering disk sources within $|b|<20^{\circ}$, weighting by X-ray flux gives an east/west ratio of $\sim 0.6$, which is opposite to earlier claims. On the other hand, if we assume that the positron annihilation rate is proportional to $L_{\mathrm{X}}^{0.5}$, where $L_{\mathrm{X}}$ is the X-ray luminosity, the east/west ratio becomes $\sim 1.4$. Here, we extracted sources whose distances were estimated in Avakyan et al. (2023). This demonstrates that understanding the luminosity dependence of positron production is crucial for interpreting spatial asymmetries. Regarding the latitude profile, LMXBs in the eastern disk region ($l=-50^{\circ}$ to $20^{\circ}$) are indeed concentrated at negative latitudes, qualitatively consistent with the southern offset in our 511 keV map. We note that, if 511 keV emission simply scales with X-ray luminosity regardless of individual LMXB properties, Sco X-1 (the brightest LMXB at $b\approx 24^{\circ}$) would produce stronger emission. These results indicate that drawing robust conclusions about LMXB contributions requires detailed modeling of positron production mechanisms in jets, escape processes from binary systems, and propagation through different ISM environments for each source, which simple estimations above cannot capture.
Figure 10: The distribution of the LMXBs in the Swift BAT 157-month survey catalog. The radius of each circle is proportional to the X-ray flux of its corresponding source.
@@ 5.2 On the hints of 511 keV emission from massive stars
$D\lesssim O(10^{26})\leavevmode\nobreak\ \mathrm{cm^{2}\leavevmode\nobreak\ s^{-1}}\leavevmode\nobreak\ .$ (11)
@@
While we have focused on massive stars as a potential source to explain the possible asymmetric bulge structure, we note that other scenarios could also explain it. For example, the chimney-like structure in the Upper Scorpius region spatially coincides with the inner regions of the Fermi and eROSITA bubbles (Su et al. 2010; Ackermann et al. 2014; Predehl et al. 2020), suggesting possible connections to past energetic outbursts from the Galactic center (e.g., Totani 2006; Crocker & Aharonian 2011). However, the inferred north-south asymmetry may not favor this scenario, since these bubble structures are symmetric about the Galactic plane. More sensitive observations, such as with COSI, are needed to test this possibility conclusively. Additionally, the asymmetry near the Galactic center may be explained by positron production through cosmic-ray interactions with asteroids (Siegert 2024). Regarding the Upper Scorpius emission, we observe hints of this emission in two different epochs (see Figure 18A), which could support either steady-state or variable emission from this region, depending on the exact exposure configuration given the variable sources within the Solar System. More detailed temporal analysis will be presented in a future work.
@@ 5.3 Implications for the future MeV gamma-ray observations
Our analysis has highlighted two key challenges in studying the 511 keV emission with INTEGRAL/SPI and interpreting the results. First, the non-uniform exposure across the sky prevents us from concluding whether some faint structures in the reconstructed image are real. As shown in Figure 1, the exposure time varies considerably, with the maximum difference in the exposure map being about three orders of magnitude. This non-uniformity can cause systematic uncertainties in the image deconvolution, making it difficult to confirm potential features. This issue is particularly problematic for the high-latitude structures we detected around ($l$, $b$) = (70, 40) and (260, -60), whose reliability remains uncertain.
@@ 6 Conclusions
Our analysis also suggests intriguing hints of new spatial features in the 511 keV emission. Notably, we find that the bulge emission may extend asymmetrically towards the northeastern direction with a faint blob. This extension is spatially coincident with the region where the enhanced 26Al emission was reported from the Scorpius–Centaurus association. This correlation could indicate a first detection of 511 keV emissions associated with 26Al from massive stars, although its significance remains marginal.
Detailed interpretation of the spatial structures is limited by two factors in the current observations. First, the non-uniform exposure across the sky may cause artifacts, such as blobs in high-latitude regions. Second, the sensitivity of current observations only allows us to detect spatial features at marginal significance beyond the bulge and disk components. Future MeV gamma-ray observations, such as with COSI, will address these issues with improved sensitivity and uniform exposure through all-sky survey capabilities. It will enable us to confirm the reported tantalizing spatial structures and reveal the origin of the positron production and its propagation in our Galaxy.
@@ Appendix D Comparison between the data and the reconstructed model
The background normalization factors are also investigated. Figure 15 shows their distribution after the image deconvolution, which is well-centered around unity. We note that one data point is present around 1.01, which appears to be an outlier. This corresponds to a pointing with a particularly short exposure time, where its orbital revolution contains only this single pointing. We also show the evolution of the background normalization factors with iteration number in Figure 16. While the background normalization for this revolution may not be fully converged, we have confirmed that excluding it from the analysis did not affect our results.
@@ Appendix E Longitude profile with different integration ranges
We show the longitude profile with different integration ranges in Figure 17. While the profile from $b=-10^{\circ}$ to $10^{\circ}$ is similar to that shown in Figure 6, the longitude profile above $b=10^{\circ}$ differs significantly from that below $b=-10^{\circ}$. Below $b=-10^{\circ}$, the profile near $l=0^{\circ}$ is nearly flat, with intensity close to the sensitivity limit. In contrast, above $b=10^{\circ}$, the profile exhibits a bright structure at $l<0^{\circ}$, which peaks at $l\sim-4^{\circ}$. Together with the plateau observed at $b=10^{\circ}$$18^{\circ}$ in the latitude profile, this suggests flux enhancement in the northeastern region. This feature also hints at a chimney-like structure in the northern region.
Figure 17: Same as Figure 5, but using different integration ranges; from $b=10$ to 2020 degrees (top), from b=10b=-10 to $10$ degrees (middle), and from b=20b=-20 to 10-10 degrees (bottom).
@@ Appendix F Image analysis using different periods
We performed image reconstruction using different periods to investigate the time variability/stability of the detected features. Figure 18 shows the reconstructed images using four independent periods (2003-2008, 2008-2013, 2013-2017, and 2017-2023) with their corresponding exposure maps. While the bulge and disk features are consistently detected in these images, some features, such as the high-latitude blobs, appeared predominantly in specific periods when exposure close to these regions was long, but they are on the edge of the deep exposure regions. It may suggest a potential systematic effect due to the combination of non-uniform exposure and the uncertainties of the off-axis response matrix. As another example, the third image shows the bright region around $(l,b)\approx(140^{\circ},40^{\circ})$, where INTEGRAL performed deep observations of $\sim 4$ Ms for the type Ia supernova SN2014J (Churazov et al. 2014; Diehl et al. 2015). We also show cumulative images starting from 2023 with increasing durations in Figure 19.
Figure 18: Reconstructed images using different independent periods of the data (from top to bottom: 2003-2008, 2008-2013, 2013-2017, and 2017-2023) with their corresponding exposure maps (right panels). In the middle panels, the maximum value is reduced to $1\times 10^{-3}$ ph cm-2 s-1 sr-1 to enhance the visibility of faint structures similar to Figure 7.
Figure 19: Same as Figure 18 but using different periods starting from 2003 with increasing durations (from top to bottom: 2003-2008, 2003-2013, 2003-2017, and the full dataset).
@@ Appendix G Bootstrap samples for individual regions
Figure 20 shows the flux distributions for individual regions derived from bootstrap samples discussed in Section 4.3.1. The blue histograms show the distributions of bootstrap samples, which were resampled from the data by adding Poisson noise. The orange histograms show those from the background-only samples. The vertical lines indicate the measured flux values from the reconstructed image shown in Figure 3.

### v3

@@ 4.3.2 Other regions
The reconstructed image shows a blob-like structure at high Galactic latitude around ($l$, $b$) = (70, 40). When defining the region as $l=60^{\circ}$ to $90^{\circ}$, $b=30^{\circ}$ to $60^{\circ}$, we measured the flux of this region to be $(7.5\pm_{\mathrm{stat}}2.4\pm_{\mathrm{sys}}3.4)\times 10^{-5}$ ph cm-2 s-1. None of the background-only bootstrap samples reproduced such a high flux in this region. Since no known bright high-energy source has been identified in this region, it could be an interesting source emitting 511 keV line predominantly, for example, a locally dense region aligned with stellar streams, Galactic winds, or transient events such as binary neutron star mergers (Fuller et al. 2019) and calcium-rich supernovae (Perets 2014), or nearby bubbles (Siegert et al. 2024).
@@
Although it is statistically significant, we need to be careful about artifacts of the image deconvolution. As seen at the bottom of Fig. 8, the exposure is not uniform in high-latitude regions. Furthermore, when splitting the full dataset into four independent periods, we found that this structure appears obviously in the image of the last period (years 2017–2023, see Appendix F and Fig. 18). From January 2022 to October 2022, SPI monitored the blazar Mkn 501, located at ($l$, $b$) = (63.6, 38.9) and apart by $\sim 6^{\circ}$ from the center of the blob-like structure. This feature may arise from systematic uncertainties in the SPI response function at the boundaries of its field of view rather than being a real astronomical source. Additionally, we observed a similar structure when performing image reconstruction in an adjacent energy band (526–532 keV) with a flux of $\sim 4\times 10^{-5}$ ph cm-2 s-1.
@@
At ($l$, $b$) = ($260$, 60-60) degrees, another structure can be seen, annotated as “South Blob” in Fig. 8. However, the non-uniform exposure could also affect this structure. To conclude the existence of these structures, uniform exposure is crucial to minimize such systematic uncertainties in image reconstruction for future 511 keV observations. In any case, this southern emission feature is too far away from the Magellanic Clouds. Interestingly, Siegert et al. (2020) saw a similar emission feature at high negative latitudes with the COSI balloon, which they also attributed to image deconvolution artifacts.
@@
Other emission features in Fig. 8, such as around the Cepheus region ($l\sim 120^{\circ}$) or at the position of M82 (in particular deep exposure of the type Ia supernova SN2014J, e.g, Diehl et al. (2014); $(l,b)\sim(140^{\circ},40^{\circ})$) appear as reasonable positron annihilation regions, but are not significant. Local features, such as Loop I or the North Polar Spur, are also not aligned with any of these serendipitous emission features. Strong individual hotspots from the filamentary structure of the Local Bubble (Siegert et al. 2024) may explain some residual emission, though not at the flux levels quoted here.
@@
Table 4: Measured fluxes and significance of the 511 keV line emission from individual regions. [^4] The chance probabilities are derived from background-only bootstrap samples.
@@ 5.1 Comparison with previous studies
For example, Weidenspointner et al. (2008) suggested that low-mass X-ray binaries with hard X-ray emissions (hard LMXBs) explain the 511 keV emission in the disk region. They reported a flux ratio of $\sim 2$ between negative and positive longitude regions, and hard LMXBs show a similar asymmetric distribution along the Galactic plane. To revisit this hypothesis using our 20-year dataset, we examined the spatial distribution of LMXBs using the Swift-BAT 157-month catalog (Lien et al. 2025). The catalog shows 71 LMXBs in the eastern hemisphere ($l<0^{\circ}$) versus 46 in the west ($l>0^{\circ}$), confirming an asymmetric distribution similar to previous findings (see Fig. 10). However, the interpretation depends critically on the assumed positron production efficiency. When considering disk sources within $|b|<20^{\circ}$, weighting by X-ray flux gives an east/west ratio of $\sim 0.6$, which is opposite to earlier claims. On the other hand, if we assume that the positron annihilation rate is proportional to $L_{\mathrm{X}}^{0.5}$, where $L_{\mathrm{X}}$ is the X-ray luminosity, the east/west ratio becomes $\sim 1.4$. Here, we extracted sources whose distances were estimated in Avakyan et al. (2023). This demonstrates that understanding the luminosity dependence of positron production is crucial for interpreting spatial asymmetries. Regarding the latitude profile, LMXBs in the eastern disk region ($l=-50^{\circ}$ to $20^{\circ}$) are indeed concentrated at negative latitudes, qualitatively consistent with the southern offset in our 511 keV map. We note that, if 511 keV emission simply scales with X-ray luminosity regardless of individual LMXB properties, Sco X-1 (the brightest LMXB at $b\approx 24^{\circ}$) would produce stronger emission. These results indicate that drawing robust conclusions about LMXB contributions requires detailed modeling of positron production mechanisms in jets, escape processes from binary systems, and propagation through different ISM environments for each source, which simple estimations above cannot capture.
Figure 10: Distribution of the LMXBs in the Swift BAT 157-month survey catalog. The radius of each circle is proportional to the X-ray flux of its corresponding source.
@@ 5.2 On the hints of 511 keV emission from massive stars
$D\lesssim O(10^{26})\penalty 10000\ \mathrm{cm^{2}\penalty 10000\ s^{-1}}\penalty 10000\ .$ (11)
@@
While we have focused on massive stars as a potential source to explain the possible asymmetric bulge structure, we note that other scenarios could also explain it. For example, the chimney-like structure in the Upper Scorpius region spatially coincides with the inner regions of the Fermi and eROSITA bubbles (Su et al. 2010; Ackermann et al. 2014; Predehl et al. 2020), suggesting possible connections to past energetic outbursts from the Galactic center (e.g., Totani 2006; Crocker & Aharonian 2011). However, the inferred north-south asymmetry may not favor this scenario, since these bubble structures are symmetric about the Galactic plane. More sensitive observations, such as with COSI, are needed to test this possibility conclusively. Additionally, the asymmetry near the Galactic center may be explained by positron production through cosmic-ray interactions with asteroids (Siegert 2024). Regarding the Upper Scorpius emission, we observe hints of this emission in two different epochs (see Fig. 18A), which could support either steady-state or variable emission from this region, depending on the exact exposure configuration given the variable sources within the Solar System. More detailed temporal analysis will be presented in a future work.
@@ 5.3 Implications for the future MeV gamma-ray observations
Our analysis has highlighted two key challenges in studying the 511 keV emission with INTEGRAL/SPI and interpreting the results. First, the non-uniform exposure across the sky prevents us from concluding whether some faint structures in the reconstructed image are real. As shown in Fig. 1, the exposure time varies considerably, with the maximum difference in the exposure map being about three orders of magnitude. This non-uniformity can cause systematic uncertainties in the image deconvolution, making it difficult to confirm potential features. This issue is particularly problematic for the high-latitude structures we detected around ($l$, $b$) = (70, 40) and (260, -60), whose reliability remains uncertain.
@@ 6 Conclusions
Our analysis also suggests intriguing hints of new spatial features in the 511 keV emission. Notably, we find that the bulge emission may extend asymmetrically toward the northeastern direction with a faint blob. This extension is spatially coincident with the region where the enhanced 26Al emission was reported from the Scorpius–Centaurus association. This correlation could indicate a first detection of 511 keV emissions associated with 26Al from massive stars, although its significance remains marginal.
Detailed interpretation of the spatial structures is limited by two factors in the current observations. First, the non-uniform exposure across the sky may cause artifacts, such as blobs in high-latitude regions. Second, the sensitivity of current observations only allows us to detect spatial features at a marginal significance beyond the bulge and disk components. Future MeV gamma-ray observations, such as with COSI, will address these issues with improved sensitivity and uniform exposure through all-sky survey capabilities. Such observations will enable us to confirm the reported tantalizing spatial structures and reveal the origin of the positron production and its propagation in our Galaxy.
@@ Appendix D Comparison between the data and the reconstructed model
The background normalization factors are also investigated. Figure 15 shows their distribution after the image deconvolution, which is well-centered around unity. We note that one data point is present around 1.01, which appears to be an outlier. This corresponds to a pointing with a particularly short exposure time, where its orbital revolution contains only this single pointing. We also show the evolution of the background normalization factors with iteration number in Fig. 16. While the background normalization for this revolution may not be fully converged, we have confirmed that excluding it from the analysis did not affect our results.
@@ Appendix E Longitude profile with different integration ranges
We show the longitude profile with different integration ranges in Fig. 17. While the profile from $b=-10^{\circ}$ to $10^{\circ}$ is similar to that shown in Fig. 6, the longitude profile above $b=10^{\circ}$ differs significantly from that below $b=-10^{\circ}$. Below $b=-10^{\circ}$, the profile near $l=0^{\circ}$ is nearly flat, with intensity close to the sensitivity limit. In contrast, above $b=10^{\circ}$, the profile exhibits a bright structure at $l<0^{\circ}$, which peaks at $l\sim-4^{\circ}$. Together with the plateau observed at $b=10^{\circ}$$18^{\circ}$ in the latitude profile, this suggests flux enhancement in the northeastern region. This feature also hints at a chimney-like structure in the northern region.
Figure 17: Same as Fig. 5, but using different integration ranges; from $b=10$ to 2020 degrees (top), from b=10b=-10 to $10$ degrees (middle), and from b=20b=-20 to 10-10 degrees (bottom).
@@ Appendix F Image analysis using different periods
We performed image reconstruction using different periods to investigate the time variability/stability of the detected features. Figure 18 shows the reconstructed images using four independent periods (2003-2008, 2008-2013, 2013-2017, and 2017-2023) with their corresponding exposure maps. While the bulge and disk features are consistently detected in these images, some features, such as the high-latitude blobs, appeared predominantly in specific periods when exposure close to these regions was long, but they are on the edge of the deep exposure regions. It may suggest a potential systematic effect due to the combination of non-uniform exposure and the uncertainties of the off-axis response matrix. As another example, the third image shows the bright region around $(l,b)\approx(140^{\circ},40^{\circ})$, where INTEGRAL performed deep observations of $\sim 4$ Ms for the type Ia supernova SN2014J (Churazov et al. 2014; Diehl et al. 2015). We also show cumulative images starting from 2023 with increasing durations in Fig. 19.
Figure 18: Reconstructed images using different independent periods of the data (from top to bottom: 2003-2008, 2008-2013, 2013-2017, and 2017-2023) with their corresponding exposure maps (right panels). In the middle panels, the maximum value is reduced to $1\times 10^{-3}$ ph cm-2 s-1 sr-1 to enhance the visibility of faint structures similar to Fig. 7.
Figure 19: Same as Fig. 18 but using different periods starting from 2003 with increasing durations (from top to bottom: 2003-2008, 2003-2013, 2003-2017, and the full dataset).
@@ Appendix G Bootstrap samples for individual regions
Figure 20 shows the flux distributions for individual regions derived from bootstrap samples discussed in Sect. 4.3.1. The blue histograms show the distributions of bootstrap samples, which were resampled from the data by adding Poisson noise. The orange histograms show those from the background-only samples. The vertical lines indicate the measured flux values from the reconstructed image shown in Fig. 3.